If an inductor and a capacitor are in series in a circuit with low resistance, what frequency may the circuit "ring" at? This is the frequency of oscillations that will naturally occur and continue once started. Give your answer in kilohertz (1000 Hz) if the inductance is 88 millihenries and the capacitance is 0.05 microfarads. Find the frequency "f", not the angular frequency omega.
Answer #1

Answer:

Given that, Capacitance C = 0.05 x 10-6 F and inductance L = 88 x 10-3 H.

Here, the Capacitor and the Inductor are in series connection with low resistance, so we can assume this as a LC series circuit. In the case of LC circuit, there is a natural resonance frequency will occur in the circuit and its expression if f = 1/2(LC)1/2

Therefore, frequency f = 1/2[(88 x 10-3 H)(0.05 x 10-6 F)]1/2 = 2400.97 H = 2.4 kHz.